OpenStudy (anonymous):

@iGreen 2. y+5=-12 a. y=-19 b. y=-17 c. y=-7 d. y=7 3. 1/4a=3 a. a=3/4 b. a=3 c. a=12 d. a=16

OpenStudy (igreen):

y + 5 = -12 Subtract 5 to both sides: -12 - 5 = -17 y = -17

OpenStudy (igreen):

1/4a = 3 Divide 1/4 to both sides or multiply by the reciprocal, 4. 4 * 3 = 12. a = 12